OpenStudy (anonymous):

You are required to choose two topics from a list of five to write about on your science test. How many different pairings are possible?

OpenStudy (anonymous):

5^2 = 25 pairings

OpenStudy (anonymous):

The answers that were given to me are this: 10 20 30 60 When I did it I also got 25.

OpenStudy (anonymous):

its C(5,2) =5*4/2=10

OpenStudy (anonymous):

Now that makes sense! Ty